OpenStudy (anonymous):

how do I simplify a^2b^3 over ab^2c

OpenStudy (anonymous):

can anybody help me please !!!

OpenStudy (the_fizicx99):

\(\ \sf \Large \dfrac{a^2b^3}{ab^2c} \) ?

OpenStudy (anonymous):

yes

OpenStudy (paki):

ab/c

OpenStudy (the_fizicx99):

\(\ \sf \dfrac{a^2}{a} = a^{2-1} = a^1 = a \) \(\ \sf \Large a = a^1 \) Same thing xD So subtract the exponents, What is \(\ \sf \dfrac{b^3}{b^2} = b^? \) I already did \(\ a \) for you :)

OpenStudy (anonymous):

b^1 right

OpenStudy (anonymous):

okay is it okay that I ask yhu nore questions like will yhu help me on them

OpenStudy (the_fizicx99):

When you divide you subtract the exponents, when you multiply you add them! Yes! \(\ \sf \dfrac{b^3}{b^2} = b^{3-2} = b^1 = b \) ! so you have \(\ \sf ab \) so far Now, there's nothing to divide C with so you're simply left with \(\ \sf \Large \dfrac{ab}{c} \)

OpenStudy (the_fizicx99):

Sure, ask :)

OpenStudy (anonymous):

simplify (-3m^5)

OpenStudy (the_fizicx99):

Are you sure it's not \(\ (-3m)^5 \) ? Otherwise that would just be \(\ -3m^5\), you can't do anything with that <--

OpenStudy (anonymous):

okay factor 5x^2+5x-30

OpenStudy (the_fizicx99):

But wait, which is it, -3m^5 or (-3m)^5 ?

OpenStudy (anonymous):

okay I got x^3-6 is that right

OpenStudy (anonymous):

(-3m^5)

OpenStudy (the_fizicx99):

Um, my internet got disconnected -_- Anyways, 5x^2+5x-30 5(x^2 + x - 6) x^2 + 3x - 2x - 6 x(x + 3) -2(x + 3) 5(x+3)(x-2)